OpenStudy (ray1998):

How do I solve this equation? 2|3x-3|=18 x=?

OpenStudy (mertsj):

1. Divide both sides by 2 2. Think of two numbers which have 9 as their absolute value 3. Set 3x-3 equal to each number. 4. Solve both equations.

OpenStudy (irishboy123):

\(|3x-3|=9\) crack on from there

OpenStudy (irishboy123):

i'd risk it and take the 3 out too

OpenStudy (gabbyalicorn):

Divide both sides by 2. 2(|3x−3|)2=182 |3x−3|=9 Solve Absolute Value. |3x−3|=9 We know either 3x−3=9or3x−3=−9 3x−3=9(Possibility 1) 3x−3+3=9+3(Add 3 to both sides) 3x=12 3x3=123(Divide both sides by 3) x=4 3x−3=−9(Possibility 2) 3x−3+3=−9+3(Add 3 to both sides) 3x=−6 3x3=−63(Divide both sides by 3) x=−2 Answer: x=4 or x=−2 :3 meow